OpenStudy (anonymous):

what is the expansion of (x+y)^6

OpenStudy (turingtest):

\[\sum_{k=0}^{6}\left(\begin{matrix}6\\k\end{matrix}\right)x^{6-k}y^k\]

OpenStudy (anonymous):

ok thats not one of the options that they give me

OpenStudy (turingtest):

do wyou know what \[\left(\begin{matrix}n\\k\end{matrix}\right)\]means?

OpenStudy (anonymous):

nope

OpenStudy (turingtest):

\[\left(\begin{matrix}n\\k\end{matrix}\right)={n!\over k!(n-k)!}\]and if that doesn't make sense to you, you can use pascals triangle to find the coefficients, or just do it all by hand (which I don't recommend)

OpenStudy (turingtest):

\[(x+y)(x+y)(x+y)(x+y)(x+y)(x+y)\]but that would be really painful

OpenStudy (anonymous):

ok know that you brought that up do you know what is the 15th entry in row 15 of pascal triangle. I have no idea how to do that

OpenStudy (turingtest):

no I don't know the trick for that, though I'm sure there is one I just derive it by hand when I use it

OpenStudy (anonymous):

oh ok well ill try to find out somewhere else for help

OpenStudy (turingtest):

the 15the row in pascals triangle should basically consist of the sequence\[a_k=\{\left(\begin{matrix}15\\k\end{matrix}\right)\}\]from k=0 to 15, so they are actually the same thing I guess... sorry I couldn't help more, just bump this question and maybe someone can better help you

OpenStudy (anonymous):

The fifteenth entry in row fifteen is \[\left(\begin{matrix}15 \\ 15\end{matrix}\right)\] \[= 15!/(15!(15 - 15)!)\] since 0! = 1, this evaluates to 1. This makes sense because the fifteenth term of the fifteenth row is the last term(1) (This assumes that you count the first row as row 1 not 0)